what is the area of equilateral triangle if the height of it is 6?
Answer by lwsshak3(11628)

what is the area of equilateral triangle if the height of it is 6?
let 2x=length of sides
Draw a perpendicular line from apex to base which cuts triangle in half and gives you a right triangle with legs 6 and x and hypotenuse=2x
By the Pythagorean Theorem:
6=√[(2x)^2-x^2)]=√[4x^2-x^2]=√3x^2=x√3
x=6/√3
2x=12/√3
Area=1/2*base*height=1/2*2x*6=1/2*12/√3*6=36/√3≈20.78 sq units